# Prompt Mixing in Diffusion Models using the Black Scholes Algorithm

### v1

@@ 4.2 Method
    Underlying stock price $S$: The underlying stock price corresponds to the current value of the asset. Let $z_{t}$ represent the predicted latents at timestep $t$. We denote $z_{0,t}$ as the latents of the final predicted image extrapolated from $z_{t}$. In other words, if the denoising process were to proceed in a straightforward manner following the same direction as computed for $z_{t}$, the latents of the final predicted image would be $z_{0,t}$. The underlying stock price, denoted as $S$, is determined by the CLIP score with respect to text prompt $P_{p}$. This CLIP score can be computed using the image decoded from $z_{0,t}$. To maintain values within the range of 00 to $100$, we multiply the CLIP score by $100$.
@@
    Strike price $K$: This represents the estimated asset price at expiry, following the last step of diffusion denoising, which predicts the final image. Based on our experience with diffusion models, where we found that a CLIP score of approximately $0.25$ indicates reasonable text-image alignment, we opted for a constant value of $0.25$.
@@
    Time to expiration $t$: $t=T-i$.
@@
    Volatility $\sigma$: We compute the volatility as the square root of the variance used by the diffusion denoising scheduler at timestep $i$.
@@
    Risk-free rate $r$: $r=1/T$.
During each iteration $i$ of diffusion denoising, we calculate the Black-Scholes score $b_{i,p}$ for each text prompt $P_{p}$, $p=1...N$ using the given variables and Equation 1 and Equation 2. Subsequently, during the next step of diffusion denoising, we condition on the text prompt with the lowest Black-Scholes score. This approach ensures that the denoising process prioritizes the text prompt associated with the minimum Black-Scholes score for generating an optimal image consistent with all prompts. Please refer to Algorithm 1 for a step-wise description of the method.
Algorithm 1 Black Scholes prompt mixing in backward diffusion enables the diffusion model generate an optimal image at the intersection of multiple text-image manifolds.
@@ 5 Experiments and Results
In line with previous research [9], we assess performance using the CLIP Score. We use two variants of the CLIP Score: (i) CLIP-combined, which computes the score by comparing the generated image against a combination of individual text prompts; and (ii) CLIP-add, which averages the CLIP scores for the generated image across each individual text prompt. To compute our metrics, we generate five images for each text prompt and use them all in the computation of our metrics.
@@
Our study examines several baselines: (i) Vanilla SD: Prompt engineering, we use the vanilla stable diffusion model and condition it on a single text prompt effectively describing all individual text prompts, (ii) Lin. Interp.: Direct combination of text embeddings, achieved through linear interpolation between text embeddings [17]. This method equally weights the text embeddings associated with each text prompt. (iii) Switching between text prompts. We consider two variations here: (iii-a) Step: Following Patashnik et. al. [27], we use one text prompt for the initial 7th to 17th denoising steps and then switch to the other text prompt for the remaining steps, (iii-b) Alt.: Following Kothandaraman et. al. [18], we alternate between the two text prompts. (iv) CLIP Min.: Score-based combination, denoted as CLIP-min, where we select the text prompt corresponding to the lowest CLIP score from the previous denoising iteration.
@@
We use the Stable Diffusion 2.1 backbone model in all our experiments. There is no training involved, we use the pretrained model to directly perform inference.
Method CLIP-combined CLIP-add
Vanilla $0.3463$ $0.3012$
Linear Int. $0.2885$ $0.2778$
Alt. Samp. $0.3445$ $0.3098$
CLIP Min. $0.3195$ $0.2955$
Step $0.322$ $0.2997$
Black Scholes $0.3469$ $0.3112$
@@
Table 1: We follow prior work [9, 3] and analyze using CLIP Scores. The Black Scholes algorithm for prompt mixing in diffusion models achieves superior results, compared to other prompt-mixing techniques, as also evidenced by the qualitative results.
5.1 Experimental Settings
@@
Object actions against backgrounds (prompt mixing w.r.t. object). We investigate the model’s ability to morph objects while considering object-level action information and scene or background context. We consider the following backgrounds, [‘walking in Times Square’, ‘skiing in Times Square’, ‘walking in a beautiful garden’, ‘surfing on the beach’, ‘eating watermelon on the beach’, ‘sitting on a sofa on the beach’, ‘watching northern lights’, ‘watching sunset at a beach’, ‘admiring the opera house in Sydney’, ‘sleeping in a cozy bedroom’, ‘admiring a beautiful waterfall in a forest’, ‘walking in a cherry blossom garden’, ‘walking in a colorful autumn forest’, ‘flying in the sky at sunset’]. The objects are [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. In total, we create $210$ prompts.
@@
Object performing action against a background (prompt mixing w.r.t. background). In this experiment, we explore how the model blends global information and backgrounds while considering objects in various scenarios. We consider the objects [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. The backgrounds and style prompts are [‘walking in Times Square’, ‘sunset’, ‘van gogh style’, ‘oil painting’, ‘a garden with tulips’, ‘northern lights’]. Overall, we have $90$ prompts in this experiment.
5.2 Analysis and Comparisons
We show qualitative results in Figure 1 and Figure 2. More qualitative results can be found in the appendix. The quantitative comparisons are in Figure 1. Our quantitative results are averaged across the metrics for each of the four sets, with $680$, $355$, $1050$, $450$ images respectively. A detailed analysis of our method and benefits over prior art is as follows:
  • Vanilla stable diffusion (SD): Vanilla stable diffusion v2.1 can create plausible combinations of the provided text prompts. However, due to hallucination issues, it is constrained by the distributions it learned during training. As a result, it struggles to capture clear characteristics of individual text prompts. For example, in Figure 1, the distinct features of the parrot, cat, dog/penguin, and sunset/penguin are missing. Similarly, in Figure 2, the avocado/raccoon, muffin, parrot, and oil painting style lack distinct characteristics. Despite this, the generated images align with the broad text prompt descriptions, resulting in a considerable CLIP score.
  • Linear interpolation [17]: Linear interpolation is ill-suited for generating an image that combines two text prompts. This limitation arises from the non-linear nature of the text-image space. While linear interpolation assumes a simple linear relationship, the actual mapping between textual descriptions and image features is intricate. Furthermore, linear interpolation can only traverse a straight path between two points in the latent space, failing to capture nuanced variations or produce novel features beyond the endpoints.
  • Alternating Sampling [18],: Alternating sampling generates features related to both text prompts in the final output image. However, the objects produced suffer from poor definition, unrealistic appearance, and low quality in many instances. Artifacts are prevalent, and several generated images appear implausible. The underlying issue lies in the algorithm’s routine alternation between the two text prompts during diffusion denoising, without adequately emphasizing the most relevant prompt at each step. Despite these limitations, the model achieves a reasonable CLIP Score by broadly aligning with the overall image description, akin to vanilla denoising.
  • Step-wise switching [18]: Step-wise switching is similarly ineffective for the same reasons as Alternating sampling.
  • CLIP-min: The CLIP-min results exhibit bias toward one of the text prompts in most cases, preventing the model from generating an image that aligns with the intersection of both prompts. This bias arises because the model overlooks the dynamics of diffusion denoising and instead selects the text prompt with the lowest CLIP score at each step, resulting in bias-related issues.
  • Black Scholes algorithm: Our approach based on the Black-Scholes algorithm effectively generates realistic images that align with the intersection of the two text prompts. These images exhibit minimal unrealistic artifacts. Additionally, our algorithm successfully preserves individual characteristics corresponding to each text prompt. By modeling the dynamics of diffusion denoising, the algorithm strategically selects the optimal text prompt at each step, achieving a balanced synthesis of both prompts. Notably, our model attains superior quantitative results.
In summary, the Black-Scholes model outperforms all previous methods for prompt mixing by generating realistic images that meticulously preserve and balance the characteristics of each individual text prompt.
@@ 5.2 Analysis and Comparisons
Figure 3: Image variations from our method using the Black Scholes model for a single prompt, starting from different random Gaussian noise initializations.
6 Conclusions, Limitations, and Future Work
In this paper, we introduce a novel perspective on prompt mixing, inspired by a probablistic prediction model and its theoretical framework from the domain of Finance. We draw parallels between the Black-Scholes algorithm used in financial markets and diffusion models. From this comparison, we derive an algorithm for prompt mixing, aiming to generate images that align with multiple text prompts. Our extensive analysis showcase both qualitative and quantitative advantages of our approach. Our study has a few limitations, which opens avenues for future work: (i) We follow prior work [9, 3] by using CLIP scores for quantitative evaluation. However, CLIP scores may not be the optimal evaluation strategy, as they can yield similar scores for images with significant quality differences. Due to its learned nature, CLIP is vulnerable to hallucination and bias issues related to its training data. Consequently, it may assign similar scores to images that exhibit general characteristics of the text prompt, without focusing on specific details and features of the scene entities described in the text prompt. To address this, future research could explore more advanced image-language models [25] for evaluation; the development of good metrics for the evaluation of generative models is a widespread issue, and could help better quantify the performance of our method. (ii) Our study focuses on prompt mixing with two prompts. Investigating the impact of using more than two prompts could be a promising direction for future work. (iii) While our method leverages diffusion models, it may not be applicable to non-Gaussian [6] diffusion models or one-step diffusion models [41], as demonstrated in recent research.
Moreover, future work on integrating our approach with recent advancements in attention guidance, layout modeling, and related areas could extend its applicability to critical downstream tasks such as image editing [17, 39, 4], compositionality [24, 3], handling complex prompts [23], text-based view synthesis [18, 19] and personalized image generation [32, 20]. While recognizing the success of recent text-to-image models like Dalle-3, Stable Diffusion 3, Firefly 3, and Imagen 3 in generating images from diverse and complex text prompts, we introduce a data-efficient approach to prompt mixing. Our method offers advantages over previous techniques for prompt-mixing, and future extension could explore integrating our model’s benefits into state-of-the-art text-to-image architectures for data efficient prompt mixing and downstream applications.
Societal impact.
While our method offers valuable tools for generative AI-based content creation, its potential misuse underscores the need for research in watermarking and deepfake detection to mitigate risks.
Acknowledgements.

### v2

@@ 4.2.1 Black-Scholes Analogy in Diffusion Models
    Underlying Stock Price, $S$: Analogous to the current value of an asset, we define $S$ as the CLIP score measuring the alignment of the generated image with a specific text prompt $P_{p}$ at the current stage. Let $z_{t}$ denote the predicted latents at timestep $t$, and $z_{0,t}$ represent the latents of the final predicted image extrapolated from $z_{t}$. The score is scaled (multiplied by 100) to lie within the range 00 to $100$.
@@
    Strike Price, $K$: This represents the asset’s price at expiration. In our formulation, $K$ is defined as the average CLIP score observed at the final diffusion step when the model generates an image using a straightforward combination of prompts. This value reflects the model’s maximum potential for that set of prompts.
@@
    Time to Expiration, $t$: In diffusion models, $t$ corresponds to the number of remaining denoising steps, i.e., $t=T-i$, where $i$ is the current step.
@@
    Volatility, $\sigma$: In finance, volatility is the standard deviation of the asset price fluctuations. Analogously, we compute $\sigma$ as the square root of the variance used by the diffusion denoising scheduler at timestep $i$, reflecting the uncertainty inherent in the diffusion process.
@@
    Risk-free Rate, $r$: In traditional finance, the risk-free rate is determined by economic factors. For our purposes, we simplify by setting $r=\frac{1}{T}$ to distribute returns equally over the entire diffusion process.
4.2.2 Black-Scholes Guided Prompt Selection
At each iteration $i$ of the diffusion denoising process, we compute the Black-Scholes score $b_{i,p}$ for each text prompt $P_{p}$ (with $p=1,\dots,N$) using the variables introduced above along with Equations 1 and 2. In the subsequent diffusion step, the model conditions on the text prompt associated with the lowest Black-Scholes score. This ensures that the denoising process prioritizes the prompt that most needs refinement, enhancing the overall alignment of the generated image with all inputs.
For a detailed step-by-step description of our method, please refer to Algorithm 1.
Algorithm 1 Black Scholes concept blending in backward diffusion enables the diffusion model generate an optimal image at the intersection of multiple text-image manifolds.
@@ 5 Experiments and Results
We evaluate performance using the following metrics:
  1. 1.
    Text-image alignment: We utilize two variants of the CLIP score:
    • CLIP-combined: This variant assesses the overall alignment with text by comparing the generated image against a combination of individual text prompts.
    • CLIP-add: This averages the CLIP scores for the generated image across each individual text prompt, reflecting alignment with each specific concept.
  2. 2.
    Preserving fidelity (or characteristics) w.r.t. each constituent concept:
    • The BLIP score is calculated by comparing the generated image to a combination of individual text prompts, measuring overall text-level alignment.
    • The DINO score evaluates the generated image against each individual text prompt, assessing how well the image preserves the characteristics of each concept. This indicates the fidelity of the attributes in the generated image relative to the individual concepts.
    Thus, while BLIP focuses on overall concept blending, DINO provides insights into the preservation of characteristics for each concept. We aim for high values in both scores, calculating a net score by multiplying the DINO and BLIP scores for each prompt and averaging across all prompts.
  3. 3.
    Generation quality: We use KID to assess the realism of the generated samples, serving as an indicator of their overall quality.
To calculate our metrics, we generate five images for each text prompt and incorporate all of them into the metric computations.
@@
Our study examines several baselines: (i) Vanilla SD: Prompt engineering, we use the vanilla stable diffusion model and condition it on a single text prompt effectively describing all individual text prompts, (ii) Linear Interpolation: Direct combination of text embeddings, achieved through linear interpolation between text embeddings [17]. This method equally weights the text embeddings associated with each text prompt. (iii) Switching between text prompts. We consider two variations here: (iii-a) Step: Following Patashnik et. al. [26], we use one text prompt for the initial 7th to 17th denoising steps and then switch to the other text prompt for the remaining steps, (iii-b) Alt.: Following Kothandaraman et. al. [18], we alternate between the two text prompts. (iv) CLIP Minimum: Score-based combination, denoted as CLIP-min, where we select the text prompt corresponding to the lowest CLIP score from the previous denoising iteration.
@@
We use the Stable Diffusion 2.1 backbone model [30] in all our experiments. There is no training involved, we use the pretrained model to directly perform inference. Our experiments are run on one NVIDIA A5000 GPU with 24 GB memory.
Method CLIP-combined ($\uparrow$) CLIP-add ($\uparrow$) BLIP $\odot$ DINO ($\uparrow$) KID ($\downarrow$) Steps ($\downarrow$) Time (s) ($\downarrow$) GPU hrs Memory (GB) ($\downarrow$)
Linear Int. [17] $0.2885$ $0.2778$ $0.2588$ $0.02851$ $50$ $6.5$ $0.001805$ $7.1$
Alt. Samp. [18] $0.3445$ $0.3098$ $0.3894$ $0.01786$ $100$ $14$ $0.00389$ $7.7$
CLIP Min. $0.3195$ $0.2955$ $0.3107$ $0.00866$ $100$ $14$ $0.00389$ $7.7$
Step [26] $0.3220$ $0.2997$ $0.3390$ $0.01709$ $100$ $14$ $0.00389$ $7.7$
Black Scholes 0.3469 0.3112 0.3912 0.01531 100 14 0.00389 7.7
@@
Table 1: We evaluate various properties using CLIP Scores, BLIP Scores, DINO, and KID. These metrics help us assess overall text alignment with the combined text prompts, the preservation of attributes related to individual concepts, and the quality of the generated images. The Black Scholes algorithm for concept blending in diffusion models achieves superior results, compared to other prompt-mixing techniques, as also evidenced by the qualitative results.
Hyperparameters.
Based on our experiments for the vanilla combination using Stable Diffusion 2.1 for the dataset under consideration, where we found that a CLIP score of approximately 0.25 indicates reasonable text-image alignment, we opted for a constant value of 0.25 for the strike price. The ordering of prompts does not matter. This is because, at every step of diffusion denoising, the algorithm chooses the prompt that should be selected by computing the Black Scholes score wrt each prompt, and this process is agnostic to the ordering of the prompts. More details on hyperparameters and the backbone architecture can be found in the appendix.
Figure 3: Our method can be extended to concept blending involving more than 2 prompts (3 in this case), as shown above.
5.1 Analysis and Comparisons
We construct a dataset with $4$ types of scenarios to analyze varying complexities, spanning:
  • Simple text prompts: A single object per prompt.
  • Multiple objects: Multiple objects described within one prompt.
  • Object actions against backgrounds: Concept blending with respect to the object.
  • Object performing action against background/style: Concept blending with respect to background and style.
More details can be found in the appendix.
Qualitative results are shown in Figure 1 and Figure 2, with additional examples provided in the appendix. Quantitative comparisons are summarized in Figure 1. A detailed analysis of our method relative to prior art is described below:
  • Vanilla Stable Diffusion (SD): Vanilla Stable Diffusion v2.1 can generate plausible combinations of the input text prompts. However, due to hallucination issues and constraints imposed by the training distribution, it struggles to capture the distinct characteristics of individual prompts. For instance, in Figure 1 the unique features of the parrot, cat, dog/penguin, and sunset/penguin are not well preserved, and similarly, Figure 2 shows a lack of distinct characteristics for avocado/raccoon, muffin, parrot, and the oil painting style.
  • Linear Interpolation [17]: Linear interpolation, which assumes a simple linear relationship, is ill-suited for blending two text prompts due to the highly non-linear nature of the text-image space. It only traverses a straight path between endpoints in the latent space, failing to capture nuanced variations or generate novel features.
  • Alternating Sampling [18]: While alternating sampling produces features related to both prompts in the final output image, it often results in poorly defined objects, unrealistic appearance, and low overall quality. Artifacts are prevalent, and many images appear implausible, leading to a high KID score. The main issue is that the method alternates between the two text prompts without emphasizing the most relevant one during each denoising step.
  • Step-wise Switching [18]: This technique suffers from similar issues as alternating sampling and fails to effectively capture the detailed characteristics of multiple text prompts.
  • CLIP-min: The CLIP-min approach tends to favor one text prompt over the other, resulting in images that do not adequately reflect the intersection of both prompts. It selects the text prompt with the lowest CLIP score at each step, thereby overlooking the dynamic nature of the diffusion denoising process.
  • Black-Scholes Algorithm: Our proposed method, based on the Black-Scholes algorithm, effectively generates realistic images that align with the intersection of the two text prompts. Our approach preserves the unique attributes of each prompt while minimizing unrealistic artifacts. By modeling the dynamics of the diffusion denoising process, our algorithm strategically selects the optimal text prompt at every step, achieving a balanced synthesis and superior quantitative performance.
In summary, our Black-Scholes model–inspired method outperforms previous prompt mixing approaches by generating realistic images that meticulously preserve and balance the features of each individual text prompt.
Computational complexity
We report the computational requirements for different methods in Table 1. All values are computed using one NVIDIA A5000 GPU, on the 2 prompts blending scenario. Our algorithm incurs minimal computational overhead over other methods, primarily arising from the computation of the CLIP score of the generated image w.r.t. all individual text prompts at every step of the diffusion denoising based generation process.
6 Conclusions, Limitations, and Future Work
In this paper, we introduced a novel approach to concept blending for text-to-image diffusion models that is deeply inspired by financial probabilistic models, particularly the Black-Scholes framework. By drawing an analogy between asset pricing and the dynamics of the diffusion process, we developed an automated algorithm that dynamically selects the most relevant text prompt at each denoising step. This approach not only enhances the alignment between the generated image and the input prompts but also offers a data-efficient alternative to conventional prompt mixing methods.
Our extensive qualitative and quantitative evaluations demonstrate that our method outperforms traditional techniques such as linear interpolation, alternating sampling, and manual prompt switching. While our approach leverages the strengths of modern diffusion models, it is important to note that it may not extend to non-Gaussian diffusion processes [6] or one-step diffusion models [39]—a limitation that opens interesting avenues for future research.
Moreover, the underlying principles of our method have significant implications beyond concept blending. They can be readily integrated into state-of-the-art text-to-image architectures, and they pave the way for critical downstream tasks including image editing [17, 37, 4], compositional synthesis [24, 3], handling complex prompts [23], text-based view synthesis [18, 19], and personalized image generation [31, 20]. Our work demonstrates that financially inspired methods have the potential to significantly enhance concept blending in text-to-image generation, thereby opening avenues for broader innovation across the field of generative AI.
Impact Statement.
Our approach advances the state-of-the-art in generative AI for content creation by harnessing the synergy between diffusion models and financial probabilistic frameworks. However, the increased realism and flexibility afforded by our method also raise significant concerns regarding potential misuse. This dual-edged potential underscores the urgent need for continued research in watermarking and deepfake detection to responsibly mitigate associated risks while enabling creative innovations.
Acknowledgements.
A.1. Experimental Settings
@@
Object actions against backgrounds (concept blending w.r.t. object). We investigate the model’s ability to morph objects while considering object-level action information and scene or background context. We consider the following backgrounds, [‘walking in Times Square’, ‘skiing in Times Square’, ‘walking in a beautiful garden’, ‘surfing on the beach’, ‘eating watermelon on the beach’, ‘sitting on a sofa on the beach’, ‘watching northern lights’, ‘watching sunset at a beach’, ‘admiring the opera house in Sydney’, ‘sleeping in a cozy bedroom’, ‘admiring a beautiful waterfall in a forest’, ‘walking in a cherry blossom garden’, ‘walking in a colorful autumn forest’, ‘flying in the sky at sunset’]. The objects are [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. In total, we create $210$ prompts.
@@
Object performing action against a background (concept blending w.r.t. background). In this experiment, we explore how the model blends global information and backgrounds while considering objects in various scenarios. We consider the objects [‘a kangaroo’, ‘a cute cat’, ‘a corgi dog’, ‘a parrot’, ‘a teddy bear’, ‘a penguin’]. The backgrounds and style prompts are [‘walking in Times Square’, ‘sunset’, ‘van gogh style’, ‘oil painting’, ‘a garden with tulips’, ‘northern lights’]. Overall, we have $90$ prompts in this experiment.
@@
Figure 4: Image variations from our method using the Black Scholes model, starting from different random Gaussian noise initializations.
Figure 5: Ablation experiment on the strike price $K$ reveals that the results are optimal at values of $K$ close to the CLIP score of the image generated in the vanilla case using a combination of the constituent text prompts.
A.1.1. Hyperparameters:
We use the Stable Diffusion 2.1 backbone in all our experiments. We use standard diffusion parameters: image size - (512, 512); classifier-free guidance scale - 7.5; DDIMScheduler with betaStart set to 0.00085, betaEnd set to 0.012, betaSchedule set to “scaledLinear”, clipSample set to False and setAlphaToOne set to False. We use $100$ inference steps in all our experiments.
For the set of prompts in our dataset under consideration, the average CLIP score for the combination of prompts in the vanilla case was $0.25$. Hence, we set the value of strike price $K$ at $0.25$.
A.2. Assumptions made by the Black Scholes Algorithm
The Black Scholes model makes five assumptions: (i) No dividends are paid out during the life of an option, (ii) Market movements are somewhat random, (iii) There are no transaction costs in buying the asset, (iv) The volatility and risk free rate of the underlying asset are known and constant, (v) The returns of the underlying asset are normally distributed. The relation between the assumptions made by the Black Scholes algorithm and diffusion models is as follows:
  • No Dividends (Assumption i): This assumption is specific to financial options and does not directly apply to diffusion models. In the context of diffusion models, we can consider that the "output" (the generated image) does not have any intermediate rewards or returns, similar to how options do not yield dividends.
  • Random Market Movements (Assumption ii): This assumption aligns well with the inherent randomness in the diffusion process. Just as stock prices are subject to unpredictable fluctuations, the generation of images in diffusion models involves stochastic processes where the final output is influenced by random noise and iterative refinement.
  • No Transaction Costs (Assumption iii): Similar to the first assumption, transaction costs are not relevant in the context of diffusion models. Instead, we can think of the computational resources and time required for generating images as analogous to transaction costs, but they do not affect the model’s fundamental operation.
  • Constant Volatility and Risk-Free Rate (Assumption iv): In diffusion models, while volatility is not explicitly defined as in financial contexts, the concept of stability in the generation process can be likened to having a consistent framework for how noise is added and refined. The “risk-free rate” does not have a direct counterpart, but the idea of a stable environment for generating images can be considered.
  • Normally Distributed Returns (Assumption v): This assumption can be related to the distribution of pixel values or features in the generated images. In many diffusion models, the outputs are conditioned on a distribution that can be approximated as normal, especially when considering the latent space representations.
In summary, while some assumptions of the Black-Scholes model are not directly applicable to diffusion models, others, particularly the randomness and distribution aspects, provide a meaningful connection, which leads to our algorithm on how concepts from finance can inform generative modeling techniques!
Figure 6: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts. For instance, it misses out on the characteristics of the dog, parrot and ice-cream. Linear interpolation generates images that are not consistent with the text prompts. CLIP Min. generates images biased towards one of the text promots. Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases. For instance, in order, the issues are: missing characteristics of dog, missing characteristics of coffee mug and artifacts in teddy bear, artifacts and missing characteristics of parrot, missing characteristics of teddy bear and artifacts in teddy bear, missing characteristics of rabbit. Our Black Scholes method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts. The text prompts are from set 1.
Figure 7: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts. For instance, it misses out on the characteristics of the dog, cat, corgi dog and basket/ teapot mixing. Linear interpolation generates images that are not consistent with the text prompts. CLIP Min. is biased towards one of the text prompts. Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases. For instance, in order, the issues are: artifacts in dog/banana and missing characteristics of dog, missing characteristics of cat, the corgi dog in the dog/apple image does not describe the dog as well as our Black Scholes method does, the dog is not well described in the fourth example, the teapot is not well described in the final image. Our method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts. The text prompts are from set 2.
Figure 8: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts. For instance, in the first and fifth image, the characteristics of the teddy bear are missing, the second, third and fourth image do not describe the cat well, Linear interpolation generates images that are not consistent with the text prompts. CLIP Min is again biased towards one of the text prompts. Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases. Specifically, the first image does not describe the subjects well, the second and third images are reasonably generated but our Black Scholes method generates a clearer image of the penguin blended cat, and teddy bear blended cat, the results of Alt. for the fourth case has artifacts and Step does not describe the cat well, the teddy bear is not well described in the fifth image. Our Black Scholes method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts. The text prompts are from set 3.
Figure 9: Vanilla stable diffusion (SD) is able to generate images satisfying both text prompts in many cases, however it misses out on the fine-grained characteristics of the individual text prompts. The first two images are cartoon-ish, and Northern lights is not clearly depicted. The third image also misses a clear description of Northern lights. In the fourth image, some of the billboards in Times Square are green, but there is no sign of Northern lights. In the fifth image, the sunset is missing. Linear interpolation generates images that are not consistent with the text prompts. CLIP Min results are biased towards one of the prompts for all images, except the second one. Alternating sampling and Step wise inference strategies generate images with a lot of artifacts, and the generated images are perceptually implausible in many cases. In the first image, the kangaroo is not generated well (almost flying in Alt’s result, and the characteristics of the kangaroo are not well generated by Step). In the second image, the characteristics of the tulip garden are not well represented. In the third image, notice how our model generates the characteristics of the sunset and Northern lights better, while ensuring that they are blended well to form a realistic image. In the fourth image, Alt and Step miss out on the characteristics of Times Square and Northern lights respectively. In the fifth image, Alt and Step miss out on the characteristics of Times Square and the sunset respectively. Our Black Scholes method is able to generate images that capture the fine-grained characteristics of objects corresponding to both text prompts, and the generated images look realistic with minimal artifacts. The text prompts are from set 4.